# First Resolution of Microlensed Images of a Binary-Lens Event

### v1

@@
National Astronomical Observatories, Chinese Academy of Science, 20A Datun Road, Chaoyang District, Beijing 100101, China
@@ A. Mérand European Southern Observatory, Karl-Schwarzschild-Straße 2, D-85748 Garching, Germany
Evelyne Alecian Univ. Grenoble Alpes, CNRS, IPAG, 38000 Grenoble, France
Carine Babusiaux Univ. Grenoble Alpes, CNRS, IPAG, 38000 Grenoble, France
@@
Maica Clavel Univ. Grenoble Alpes, CNRS, IPAG, 38000 Grenoble, France
@@
Mathis Houlle Université Còte d’Azur, Observatoire de la Còte d’Azur, CNRS, Laboratoire Lagrange, Nice, France
@@
Olivier Lai Université Còte d’Azur, Observatoire de la Còte d’Azur, CNRS, Laboratoire Lagrange, Nice, France
@@
James Leftley Université Còte d’Azur, Observatoire de la Còte d’Azur, CNRS, Laboratoire Lagrange, Nice, France
Bruno Lopez Université Còte d’Azur, Observatoire de la Còte d’Azur, CNRS, Laboratoire Lagrange, Nice, France
@@
Florentin Millour Université Còte d’Azur, Observatoire de la Còte d’Azur, CNRS, Laboratoire Lagrange, Nice, France
@@
Romain Petrov Université Còte d’Azur, Observatoire de la Còte d’Azur, CNRS, Laboratoire Lagrange, Nice, France
@@
Sylvie Robbe-Dubois Université Còte d’Azur, Observatoire de la Còte d’Azur, CNRS, Laboratoire Lagrange, Nice, France
@@
Jules Scigliuto Université Còte d’Azur, Observatoire de la Còte d’Azur, CNRS, Laboratoire Lagrange, Nice, France
@@ Abstract
We resolve the multiple images of the binary-lens microlensing event ASASSN-22av using the GRAVITY instrument of the Very Large Telescope Interferometer (VLTI). The light curves show weak binary perturbations, complicating the analysis, but the joint modeling with the VLTI data breaks several degeneracies, arriving at a strongly favored solution. Thanks to precise measurements of angular Einstein radius $\theta_{\rm E}=0.726\pm 0.002$ mas and microlens parallax, we determine that the lens system consists of two M dwarfs with masses of $M_{1}=0.261\pm{0.009}\,M_{\odot}$ and $M_{2}=0.252\pm{0.017}\,M_{\odot}$, a projected separation of $r_{\perp}=7.42\pm 0.33\,{\rm AU}$ and a distance of $D_{\rm L}=2.31\pm 0.09$ kpc. The successful VLTI observations of ASASSN-22av open up a new path for studying intermediate-separation (i.e., a few AUs) stellar-mass binaries, including those containing dark compact objects such as neutron stars and stellar-mass black holes.
@@ 1 Introduction
In recent years, the enhanced sensitivity of advanced instrumentation (Eisenhauer et al., 2023) at the Very Large Telescope Interferometer (VLTI) of the European Southern Observatory (ESO) has facilitated interferometric resolution of microlensed images (Delplancke et al., 2001), with the first successful case achieved by Dong et al. (2019) using the GRAVITY instrument (GRAVITY Collaboration et al., 2017). By enabling direct mass determinations, interferometric microlensing opens up a new path into studying Galactic stellar-mass systems, including dark compact objects such as neutron stars and black holes.
The observable effect of microlensing (Einstein, 1936; Paczyński, 1986) is the magnified flux of a background star (i.e., the “source”) due to the gravitational deflection by a foreground object (i.e., the “lens”), and thus microlensing is a sensitive probe to the mass of the lens system, irrespective of its brightness. However, the most easily extracted quantity, the Einstein timescale $t_{\rm E}$ is a degenerate combination of the lens mass $M_{\rm L}$, the lens-source relative parallax $\pi_{\rm rel}$ and the proper motion $\mu_{\rm rel}$,
@@
where $\theta_{\rm E}$ is the angular Einstein radius, which is on the order of milliarcseconds for stellar-mass microlenses in the Galaxy. To break this degeneracy, a common approach is to measure both $\theta_{\rm E}$ and the so-called “microlens parallax” ${\bm{\pi}}_{\rm E}$ to determine the lens mass (Gould, 2000, 2004),
@@
where ${\bm{\pi}}_{\rm E}$ is a two-dimensional vector (Gould, 2004) along the same direction as the lens-source relative proper motion vector ${\bm{\mu}}_{\rm rel}$,
@@
The key advantage of the interferometric microlensing method is that, it yields not only precise $\theta_{\rm E}$ from measuring the angular separation of images, but also constraints on the direction of ${\bm{\pi}}_{\rm E}$.
One often obtains “one-dimensional” microlens parallax from the light curves, that is, the component of ${\bm{\pi}}_{\rm E}$ parallel to the Earth’s acceleration is generally much better constrained than the perpendicular component (Gould et al., 1994; Smith et al., 2003; Gould, 2004). As demonstrated by Dong et al. (2019), Zang et al. (2020), and Cassan et al. (2022), VLTI microlensing can drastically improve the ${\bm{\pi}}_{\rm E}$ measurements from accurate directional constraints via measuring the position angles of images.
A long-adopted method to determine $\theta_{\rm E}$ is via the “finite-source effects” (Gould, 1994a), when the source transits/approaches a caustic (a set of positions that, for a point source, induce infinite magnification). Modeling the finite-source effects in the microlensing light curve measures the ratio between $\theta_{\rm E}$ and the angular radius of the source star $\theta_{*}$, where $\theta_{*}$ can be derived from photometric properties of the source. For a single lens, the caustic is point-like, so finite-source detections in stellar single-lens events are usually restricted to rare cases reaching high peak magnification of $A_{\rm peak}\gtrsim 100$. In contrast, binary lenses have extended caustic structures, and thus, finite-source effects manifest more often in binary-lens events. The first microlens mass measurement by An et al. (2002) is from such a caustic-crossing binary-lens event. However, this method is ineffective for studying events with weak binary perturbations (e.g., non-caustic-crossing). One such weak-binary event was OGLE-2005-SMC-001, for which Dong et al. (2007) detected the space-based microlens parallax (Refsdal, 1966; Gould, 1994b) for the first time, and since no finite-source effects were detected, it was not possible to definitively determine the lens mass due to the absence of a $\theta_{\rm E}$ measurement. While the exact fraction of weak binary events is still an open question, Zhu et al. (2014) predicted and Jung et al. (2022) confirmed that half of planetary events would lack caustic crossings despite their continuous and high-cadence coverages, implying that weak binaries may encompass a significant fraction of binary-lens parameter space.
@@
The VLTI-microlensing observations published so far have been exclusively on single-lens events (or hosts of extremely low-mass companions). Here, we analyze ASASSN-22av, the first binary-lens microlensing event with multiple images resolved by VLTI.
@@ 2 Observations and Data Reduction
We began to conduct VLTI GRAVITY observations in the night of UT 2022-01-21, and subsequently made ongoing observations until 2022-01-25. We discuss the VLTI observations and data reduction in § 2.1. We also started photometric follow-up observations on UT 2022-01-20 (see § 2.2 for detailed descriptions of photometric data). On UT 2022-01-21, our light-curve modeling suggested significant deviations from the best-fit PSPL model, and we found that a finite-source point-lens (FSPL) model was now preferred. Aiming at performing “stellar atmosphere tomography” (see, e.g., Albrow et al., 2001; Afonso et al., 2001, and references therein) near the caustic exit, we initiated a campaign of time-series high-resolution spectroscopy. On UT 2022-01-24, we noticed that the light curve declined faster than expected from the best-fit FSPL model and considered the possibility of a binary-lens perturbation, which was confirmed by our later analysis (see § 3 for detailed discussions on light-curve modeling). We obtained high-resolution spectra from the High Resolution Spectrograph (HRS) on the Southern African Large Telescope (SALT; Buckley et al. 2006), the MIKE spectrograph (Bernstein et al., 2003) on the Magellan Clay 6.5 m telescope, the High Efficiency and Resolution Multi Element Spectrograph (HERMES; Sheinis et al. 2015) on the Anglo-Australian Telescope (AAT), and the HARPS spectrograph (Mayor et al., 2003) on the ESO’s 3.6 m telescope from UT 2022-01-25 to UT 2022-01-30. The full description of the spectroscopic data and the stellar atmosphere tomography analysis will be presented in a future paper. In this work, we utilize the stellar parameters of the source extracted from the HERMES data presented in § 2.3.
@@ 2.1 Interferometric Data
We observed ASASSN-22av with the VLTI GRAVITY instrument (GRAVITY Collaboration et al., 2017) on the four nights of 21, 22, 24, and 25 January 2022. The observations were executed with the single-field on-axis mode at medium-resolution ($R=\lambda/\Delta\lambda\simeq 500)$. We use all data except those on 2022/01/21 when there was significant flux loss in UT2 and poor fringe-tracking performance. The remaining three nights of data are reduced with the standard GRAVITY pipeline (version 1.6.6), and there are 2, 4, and 2 exposures in the nights of 22, 24 and 25, respectively. Throughout this paper, we refer to these three nights as Nights 1, 2, and 3, respectively. We first use the Python script run_gravi_reduced.py to reduce the raw data up to the application of the pixel-to-visibility matrix (P2VM). The default options are used except for adopting --gravity_vis.output-phase-sc=SELF_VISPHI to calculate the internal differential phase between each spectral channel. The pipeline performed the bias and sky subtraction, flat fielding, wavelength calibration, and spectral extraction. Application of the P2VM converts the pixel detector counts into complex visibilities, taking into account all instrumental effects, including relative throughput, coherence, phase shift, and cross-talk. The dark, bad pixel, flat field, wavelength calibration, and P2VM matrix data are reduced from the daily calibration data obtained close in time to our observations. We then use run_gravi_trend.py to calibrate the closure phase with the calibrator observation of HD134122 next to the science target observations.
The observations employed 4 UTs and thus there are four closure phases for every spectral channel. In principle, the four closure phases are not independent and should sum up to zero. We examine the calibrator data and find that the standard deviation of the sum of the closure phases is $\sim 0.15^{\circ}$, comparable to the pipeline-reported closure phase uncertainties in the science data. Thus, we take the simplified approach and treat the four closure phases as uncorrelated observables rather than the more sophisticated approach of treating the correlated noise employed by Kammerer et al. (2020). On each night, we have two exposures for the calibrator star, and we find the closure phase differences of the two exposures have scatters of $0.21^{\circ},0.25^{\circ},0.35^{\circ}$ on Night 1, 2 and 3, respectively. We adopt these values as the uncertainties introduced in the calibration process and add them quadratically to the closure-phase errors.
The medium-resolution spectrograph samples 233 wavelengths spanning $2.0-2.4\,\rm um$. To speed up computation, the 233 closure phases are binned into 10 phases in our modeling process. The first and last phases are discarded due to the low signal-to-noise ratio and significant deviation from the model. See Appendix A for the description of the binning process.
@@ 2.2 Photometric Data
We obtained follow-up imaging data taken with 1-m telescopes of Las Cumbres Observatory Global Telescope Network (LCOGT; Brown et al., 2013) in the Sloan $g^{\prime}r^{\prime}i^{\prime}$ filters, the 0.4-m telescope of Auckland Observatory (AO) at Auckland (New Zealand) in the $g^{\prime}r^{\prime}i^{\prime}$ filters, and the 30-cm Perth Exoplanet Survey Telescope (PEST) in the $g^{\prime}r^{\prime}i^{\prime}$ filters.
@@
We use the PmPyeasy pipeline (Chen et al., 2022) to do the photometry on these follow-up images. For LCOGT and PEST data, we conduct image subtractions with the hotpants [^3] https://github.com/acbecker/hotpants code (Becker, 2015), and then we employ aperture photometry on the subtracted images. For the AO data, we performed point-spread-function (PSF) photometry using DoPHOT (Schechter et al., 1993).
@@
We co-added and analyzed the $g^{\prime}$-band ASAS-SN images using the standard ASAS-SN photometric pipeline based on the ISIS image-subtraction code (Alard & Lupton, 1998; Alard, 2000). The baseline ($g^{\prime}\sim 15.5$) is too faint to obtain precise photometry for ASAS-SN, and we analyze the magnified ASAS-SN epochs with ${\rm HJD^{\prime}=HJD}-2450000\in(9577.857,9677.608)$. There are data from three ASAS-SN cameras (be, bi and bm), and they are analyzed independently in the modeling.
@@
This event was announced [^4] https://gsaweb.ast.cam.ac.uk/alerts/alert/Gaia22ahy/ as Gaia22ahy by Gaia Science Alerts (GSA; Hodgkin et al., 2021) on UT 2022-01-25.4. The Gaia $G$-band light curve available on GSA website does not have uncertainties, and we adopt the same error bar ($\sigma_{G}=0.02$) for all data points. We noticed that the data in 2015 exhibit systematically larger scatter than the rest of the Gaia baseline. By inspecting a few other GSA microlensing light curves (Wyrzykowski et al., 2023), we found similar patterns at the beginning of the Gaia mission, possibly due to the continued contamination by water ice (Gaia Collaboration et al., 2016). We conservatively remove all Gaia data before the last decontamination run on 2016/08/22 (Riello et al., 2021) from our analysis.
We follow the standard procedure (Yee et al., 2012) of re-normalizing the error bars such that $\chi^{2}/{\rm d.o.f}$ is unity for each data set for the best-fit model. We reject the 3 $\sigma$ outliers and thereby exclude 4 out of 983 photometric data points from the analysis (the rejected data consist of one data point each from LCOGT $r^{\prime}$ and $i^{\prime}$, and one epoch each from the ASAS-SN bi and bm cameras). Figure 1 shows the multi-band light curves of ASASSN-22av.
2.3 HERMES Spectroscopic Data & Stellar Parameters

### v3

@@
National Astronomical Observatories, Chinese Academy of Science, 20A Datun Road, Chaoyang District, Beijing 100101, People’s Republic of China
@@ A. Mérand European Southern Observatory, Karl-Schwarzschild-Straße 2, D-85748 Garching, Germany
Mathis Houlle Université Còte d’Azur, Observatoire de la Còte d’Azur, CNRS, Laboratoire Lagrange, CS 34229, F-06304 Nice cedex 4, France
@@
Olivier Lai Université Còte d’Azur, Observatoire de la Còte d’Azur, CNRS, Laboratoire Lagrange, CS 34229, F-06304 Nice cedex 4, France
@@
James Leftley Université Còte d’Azur, Observatoire de la Còte d’Azur, CNRS, Laboratoire Lagrange, CS 34229, F-06304 Nice cedex 4, France
Bruno Lopez Université Còte d’Azur, Observatoire de la Còte d’Azur, CNRS, Laboratoire Lagrange, CS 34229, F-06304 Nice cedex 4, France
@@
Florentin Millour Université Còte d’Azur, Observatoire de la Còte d’Azur, CNRS, Laboratoire Lagrange, CS 34229, F-06304 Nice cedex 4, France
@@
Romain Petrov Université Còte d’Azur, Observatoire de la Còte d’Azur, CNRS, Laboratoire Lagrange, CS 34229, F-06304 Nice cedex 4, France
@@
Sylvie Robbe-Dubois Université Còte d’Azur, Observatoire de la Còte d’Azur, CNRS, Laboratoire Lagrange, CS 34229, F-06304 Nice cedex 4, France
@@
Jules Scigliuto Université Còte d’Azur, Observatoire de la Còte d’Azur, CNRS, Laboratoire Lagrange, CS 34229, F-06304 Nice cedex 4, France
@@ Abstract
We resolve the multiple images of the binary-lens microlensing event ASASSN-22av using the GRAVITY instrument of the Very Large Telescope Interferometer (VLTI). The light curves show weak binary-lens perturbations, complicating the analysis, but the joint modeling with the VLTI data breaks several degeneracies, arriving at a strongly favored solution. Thanks to precise measurements of angular Einstein radius $\theta_{\rm E}=0.724\pm 0.002$ mas and microlens parallax, we determine that the lens system consists of two M dwarfs with masses of $M_{1}=0.258\pm{0.008}\,M_{\odot}$ and $M_{2}=0.130\pm{0.007}\,M_{\odot}$, a projected separation of $r_{\perp}=6.83\pm 0.31\,{\rm au}$ and a distance of $D_{\rm L}=2.29\pm 0.08$ kpc. The successful VLTI observations of ASASSN-22av open up a new path for studying intermediate-separation (i.e., a few astronomical units) stellar-mass binaries, including those containing dark compact objects such as neutron stars and stellar-mass black holes.
@@
Gravitational microlensing (672); Binary lens microlensing(2136); Optical interferometry (1168)
@@ 1 Introduction
In recent years, the enhanced sensitivity of advanced instrumentation (Eisenhauer et al., 2023) at the Very Large Telescope Interferometer (VLTI) of the European Southern Observatory (ESO) has facilitated the interferometric resolution of microlensed images (Delplancke et al., 2001), with the first successful case achieved by Dong et al. (2019) using the GRAVITY instrument (GRAVITY Collaboration et al., 2017). By enabling direct mass determinations, interferometric microlensing opens up a new path into studying Galactic stellar-mass systems, including dark compact objects such as neutron stars and black holes.
The observable effect of microlensing (Einstein, 1936; Paczyński, 1986) is the magnified flux of a background star (i.e., the “source”) due to the gravitational deflection by a foreground object (i.e., the “lens”), and thus microlensing is a sensitive probe of the mass of the lens system, irrespective of its brightness. However, the most easily extracted quantity, the Einstein timescale $t_{\rm E}$ is a degenerate combination of the lens mass $M_{\rm L}$, the lens-source relative parallax $\pi_{\rm rel}$ and the proper motion $\mu_{\rm rel}$:
@@
where $\theta_{\rm E}$ is the angular Einstein radius, which is of the order of milliarcseconds for stellar-mass microlenses in the Galaxy. To break this degeneracy, a common approach is to measure both $\theta_{\rm E}$ and the so-called “microlens parallax” ${\bm{\pi}}_{\rm E}$ to determine the lens mass (Gould, 2000, 2004):
@@
where ${\bm{\pi}}_{\rm E}$ is a two-dimensional vector (Gould, 2004) along the same direction as the lens-source relative proper motion vector ${\bm{\mu}}_{\rm rel}$:
@@
The key advantage of the interferometric microlensing method is that it yields not only precise $\theta_{\rm E}$, from measuring the angular separation of images, but also constraints on the direction of ${\bm{\pi}}_{\rm E}$.
One often obtains “one-dimensional” microlens parallax from the light curves; that is, the component of ${\bm{\pi}}_{\rm E}$ parallel to the Earth’s acceleration is generally much better constrained than the perpendicular component (Gould et al., 1994; Smith et al., 2003; Gould, 2004). As demonstrated by Dong et al. (2019), Zang et al. (2020), and Cassan et al. (2022), VLTI microlensing can drastically improve the ${\bm{\pi}}_{\rm E}$ measurements from accurate directional constraints via measuring the position angles of images.
A long-adopted method for determining $\theta_{\rm E}$ is via the “finite-source effects” (Gould, 1994a), when the source transits/approaches a caustic (a set of positions that, for a point source, induce infinite magnification). Modeling the finite-source effects in the microlensing light curve measures the ratio between $\theta_{\rm E}$ and the angular radius of the source star $\theta_{*}$, where $\theta_{*}$ can be derived from the photometric properties of the source. For a single lens, the caustic is point-like, so finite-source detections in stellar single-lens events are usually restricted to rare cases reaching a high peak magnification of $A_{\rm peak}\gtrsim 100$. In contrast, binary lenses have extended caustic structures, and thus finite-source effects manifest more often in binary-lens events. The first microlens mass measurement by An et al. (2002) is from such a caustic-crossing binary-lens event. However, this method is ineffective for studying events with weak binary-lens perturbations (e.g., non-caustic-crossing). One such weak-binary event was OGLE-2005-SMC-001, for which Dong et al. (2007) detected the space-based microlens parallax (Refsdal, 1966; Gould, 1994b) for the first time, and since no finite-source effects were detected, it was not possible to definitively determine the lens mass due to the absence of a $\theta_{\rm E}$ measurement. While the exact fraction of weak binary-lens events is still an open question, Zhu et al. (2014) predicted and Jung et al. (2022) confirmed that half of planetary events would lack caustic crossings, despite their continuous and high-cadence coverages, implying that weak binaries may encompass a significant fraction of binary-lens parameter space.
@@
The VLTI microlensing observations published so far have been exclusively on single-lens events (or hosts of extremely low-mass companions). Here, we analyze ASASSN-22av, the first binary-lens microlensing event with multiple images resolved by VLTI.
@@ 2 Observations and Data Reduction
We began to conduct VLTI GRAVITY observations on the night of UT 2022 January 21 and subsequently made ongoing observations until 2022 January 25. We discuss the VLTI observations and data reduction in § 2.1. We also started photometric follow-up observations on UT 2022 January 20 (see § 2.2 for detailed descriptions of photometric data). On UT 2022 January 21, our light-curve modeling suggested significant deviations from the best-fit PSPL model, and we found that a finite-source point-lens (FSPL) model was now preferred. Aiming at performing “stellar atmosphere tomography” (see, e.g., Albrow et al., 2001; Afonso et al., 2001, and references therein) near the caustic exit, we initiated a campaign of time-series high-resolution spectroscopy. On UT 2022 January 24, we noticed that the light curve declined faster than expected from the best-fit FSPL model and considered the possibility of a binary-lens perturbation, which was confirmed by our later analysis (see § 3 for detailed discussions on light-curve modeling). We obtained high-resolution spectra from the High Resolution Spectrograph (HRS) on the Southern African Large Telescope (SALT; Buckley et al. 2006), the MIKE spectrograph (Bernstein et al., 2003) on the Magellan Clay 6.5 m telescope, the High Efficiency and Resolution Multi Element Spectrograph (HERMES; Sheinis et al. 2015) on the Anglo-Australian Telescope (AAT), and the HARPS spectrograph (Mayor et al., 2003) on the ESO’s 3.6 m telescope, from UT 2022 January 25 to UT 2022 January 30. The full description of the spectroscopic data and the stellar atmosphere tomography analysis will be presented in a future paper. In this work, we utilize the stellar parameters of the source extracted from the HERMES data presented in § 2.3.
@@ 2.1 Interferometric Data
We observed ASASSN-22av with the GRAVITY instrument (GRAVITY Collaboration et al., 2017) and four 8 m Unit Telescopes (UTs) on the four nights of 2022 January 21, 22, 24, and 25. The observations were executed with the single-field on-axis mode at medium resolution ($R=\lambda/\Delta\lambda\simeq 500)$. We use all data except those on 2022 January 21, when there was significant flux loss in UT2 and poor fringe-tracking performance. The remaining three nights of data are reduced with the standard GRAVITY pipeline (version 1.6.6), and there are two, four, and two exposures on the nights of 22, 24, and 25, respectively. Throughout this paper, we refer to these three nights as Nights 1, 2, and 3, respectively. We first use the Python script run_gravi_reduced.py to reduce the raw data up to the application of the pixel-to-visibility matrix (P2VM). The default options are used except for adopting --gravity_vis.output-phase-sc=SELF_VISPHI to calculate the internal differential phase between each spectral channel. The pipeline performed the bias and sky subtraction, flat-fielding, wavelength calibration, and spectral extraction. Application of the P2VM converts the pixel detector counts into complex visibilities, taking into account all instrumental effects, including relative throughput, coherence, phase shift, and crosstalk. The dark, bad-pixel, flat-field, wavelength calibration, and P2VM matrix data are reduced from the daily calibration data obtained close in time to our observations. We then use run_gravi_trend.py to calibrate the closure phase with the calibrator observation of HD 134122 next to the science target observations.
The observations employed four UTs and thus there are four closure phases for every spectral channel. In principle, the four closure phases are not independent and should sum up to zero. We examine the calibrator data and find that the standard deviation of the sum of the closure phases is $\sim 0.^{\circ}15$, comparable to the pipeline-reported closure-phase uncertainties in the science data. Thus, we take the simplified approach and treat the four closure phases as uncorrelated observables rather than the more sophisticated approach of treating the correlated noise employed by Kammerer et al. (2020). On each night, we have two exposures for the calibrator star, and we find the closure-phase differences of the two exposures have scatters of $0.^{\circ}21,0.^{\circ}25,0.^{\circ}35$ on Night 1, 2, and 3, respectively. We adopt these values as the uncertainties introduced in the calibration process and add them quadratically to the closure-phase errors.
The medium-resolution spectrograph samples 233 wavelengths spanning $2.0-2.4\,\rm\mu m$. To speed up computation, the 233 closure phases are binned into 10 phases in our modeling process. The first and last phases are discarded due to the low signal-to-noise ratios and significant deviations from the model. See Appendix A for the description of the binning process.
@@ 2.2 Photometric Data
We obtained follow-up imaging data taken with the 1 m telescopes of the Las Cumbres Observatory Global Telescope Network (LCOGT; Brown et al., 2013) in the Sloan $g^{\prime}r^{\prime}i^{\prime}$ filters, the 0.4 m telescope of Auckland Observatory (AO) at Auckland (New Zealand) in the $g^{\prime}r^{\prime}i^{\prime}$ filters, and the 30 cm Perth Exoplanet Survey Telescope (PEST) in the $g^{\prime}r^{\prime}i^{\prime}$ filters.
@@
We use the PmPyeasy pipeline (Chen et al., 2022) to do the photometry on these follow-up images. For LCOGT and PEST data, we conduct image subtractions with the hotpants [^3] https://github.com/acbecker/hotpants code (Becker, 2015), and then we employ aperture photometry on the subtracted images. For the AO data, we performed point-spread function (PSF) photometry using DoPHOT (Schechter et al., 1993).
@@
We coadded and analyzed the $g^{\prime}$-band ASAS-SN images using the standard ASAS-SN photometric pipeline based on the ISIS image-subtraction code (Alard & Lupton, 1998; Alard, 2000). The baseline ($g^{\prime}\sim 15.5$) is too faint to obtain precise photometry for ASAS-SN, and we analyze the magnified ASAS-SN epochs with ${\rm HJD^{\prime}=HJD}-2450000\in(9577.857,9677.608)$. There are data from three ASAS-SN cameras (be, bi and bm), and they are analyzed independently in the modeling.
@@
This event was announced [^4] https://gsaweb.ast.cam.ac.uk/alerts/alert/Gaia22ahy/ as Gaia22ahy by Gaia Science Alerts (GSA; Hodgkin et al., 2021) on UT 2022-01-25.4. The Gaia $G$-band light curve available on GSA website does not have uncertainties, and we adopt the same error bar ($\sigma_{G}=0.02$) for all data points. We noticed that the data in 2015 exhibit systematically larger scatter than the rest of the Gaia baseline. By inspecting a few other GSA microlensing light curves (Wyrzykowski et al., 2023), we found similar patterns at the beginning of the Gaia mission, possibly due to the continued contamination by water ice (Gaia Collaboration et al., 2016). We conservatively remove all Gaia data before the last decontamination run on 2016 August 22 (Riello et al., 2021) from our analysis.
[…]
2.3 HERMES Spectroscopic Data and Stellar Parameters